# Experimental Demonstration of a Quantum-Optimal Coronagraph Using Spatial Mode Sorters

### v1

@@ C.1 Cross-Talk Matrix $\Omega$
  • $Z\in\mathbb{R}^{K\times P}$: Zernike expansion coefficients for each off-axis source location.
  • $M\in\mathbb{R}^{K\times P}$: Measurements of intensity in each mode channel.
  • $\Gamma_{\Omega}=\nabla_{\Omega}\mathcal{F}$: Gradient of the objective function with respect to the cross-talk matrix.
  • $G(\Omega)=\Gamma_{\Omega}\Omega^{\dagger}-\Omega\Gamma^{\dagger}_{\Omega}$: Riemannian gradient of the objective function (Hermitian).
The number of off-axis point-source locations measured is $P=|\mathcal{Y}|$. We solve this optimization using the constrained gradient optimization algorithm detailed in Abrudan et al. (2005). The algorithm is briefly summarized as follows. Instantiate $\Omega_{0}=I$ with learning rate $\mu>0$. Then, at iteration $t$, we update the cross-talk matrix as
@@
where $G_{t}=G(\Omega_{t})$. The iterations are performed until convergence. In Figure 4a, we show the measured mode intensity data alongside the theoretical mode intensities after applying the best-fit cross-talk matrix shown in 4b.
C.2 Dark Noise Rate $\lambda_{D}$ and Structured Background $\mathbf{s}$
@@
C.3 Signal Photon Rate $\lambda_{0}$
@@
$r_{j}^{(i)}(\Lambda_{0},\vec{r}_{e})=Y_{j}^{(i)}(\vec{r}_{e})-\Lambda_{0}p_{j}(\vec{r}_{e})-\Lambda_{B}p_{B_{j}}$
@@
Figure 10: Histogram over estimates of the scene photon emission rate $\lambda_{0}$ from which we compute the best linear unbiased estimator (dashed red) $\check{\lambda}_{0}=3.17\times 10^{6}$ [Photons $\cdot T^{-1}$]. The edges of the FWHM are used in Figure 6 to determine the range of Cramer-Rao Bounds associated with our experiment.
C.4 Throughput and Bias-Scaling
[…]
@@ C.4 Throughput and Bias-Scaling
Appendix D Function Vectorization
Let $\Delta x,\Delta y$ be the horizontal and vertical pixel pitch dimensions for specific camera, and let $\{(x_{i},y_{i})\}_{i=1}^{M}$ be the set of points corresponding to the center of each pixel. The $\text{vec}(\cdot)$ operation takes a well-behaved function $f(x,y)$ over the plane and converts it to a column vector $\mathbf{f}$ whose entries are given by,
[…]
Appendix E Numerical Comparisons to Other Coronagraphs
Figure 12: Comparison of the Cramer-Rao Lower Bounds on imprecision of an unbiased estimator of the exoplanet position for different coronagraphs operating under the same empirical background model found in our experimental system. The quantum limit curve (black dashed) is for a shot-noise-limited situation without any background contribution.
@@ Appendix E Numerical Comparisons to Other Coronagraphs
$\mathbf{q}(\vec{r}_{0})=|U\Sigma V^{\dagger}\mathbf{\psi}(\vec{r}_{0})|^{2}$ (21)
where $U,V\in\mathbb{C}^{M\times M}$ are unitary matrices and $\Sigma\in\mathbb{R}^{M\times M}$ is a diagonal matrix of singular values. Figure 12 shows the Cramer-Rao lower bounds associated with each coronagraph. We see that the idealized version of our experimental setup (Experiment 4-Mode) outperforms both the PIAACMC and Vortex coronagraph for over sub-diffraction exoplanet locations in the range $|y_{e}|<=0.4\sigma$ which corresponds to where the CFI of the truncated mode basis begins to depart from the QFI limit as shown in Figure 3b.

### v3

@@ D.1 Cross-Talk Matrix $\Omega$
  • $V\in\mathbb{R}^{K\times P}$: Measurements of intensity in each mode channel. The $i^{th}$ column of $V$ is the measurement vector $\mathbf{v}_{i}$ collected for the $i^{th}$ position of the point source.
  • $\tilde{V}(\Omega)=(\Omega^{\dagger}Z)\odot(\Omega^{\dagger}Z)^{*}\in\mathbb{R}^{K\times P}:$ Forward model of the expected intensities measured in each mode channel under the unitary cross-talk matrix $\Omega$.
  • $Z\in\mathbb{R}^{K\times P}$: Mode expansion coefficients for each off-axis source location in the calibration scan. The $i^{th}$ column of $Z$ is the vector $\mathbf{z}_{i}=\Psi^{\dagger}\mathbf{u}(y_{i})\in\mathbb{C}^{K}$ for $y_{i}\in\mathcal{Y}$.
[…]
    $\bigg{(}\frac{\partial\mathcal{F}}{\partial\Omega^{*}}\bigg{)}(\Omega)=-2Z\,[(V-\tilde{V}(\Omega))\odot(\Omega^{\dagger}Z)]^{\dagger}$ (19)
  • $G(\Omega)=\Gamma_{\Omega}\Omega^{\dagger}-\Omega\Gamma^{\dagger}_{\Omega}$: Riemannian gradient of the objective function over the Lie group $U(K)$.
We solve this optimization using the constrained descent algorithm detailed in Abrudan et al. (2008, 2005). The algorithm is briefly summarized as follows. Instantiate $\Omega_{0}=I$ with learning rate $\mu>0$. Then, at iteration $t$, we update the cross-talk matrix as,
@@
where $G_{t}=G(\Omega_{t})$ and $\texttt{expm}(\cdot)$ is a numerical matrix exponentiation function. The iterations are performed until convergence. In Figure 4(a), we show the measured mode intensity data alongside the theoretical mode intensities after applying the best-fit cross-talk matrix shown in 5.
D.2 Dark Noise Rate $\lambda_{D}$ and Structured Background $\mathbf{s}$
@@
D.3 Signal Photon Rate $\lambda_{0}$
@@
$r_{j}^{(i)}(\Lambda_{0},\vec{r}_{e})=Y_{j}^{(i)}(\vec{r}_{e})-\Lambda_{0}p_{j}(\vec{r}_{e})-\Lambda_{B}p_{B_{j}}$ (25)
@@
Figure 12: Histogram over estimates of the scene photon emission rate $\lambda_{0}$ from which we compute the best linear unbiased estimator (dashed red) $\check{\lambda}_{0}=3.17\times 10^{6}$ [Photons $\cdot T^{-1}$]. The edges of the FWHM are used in Figure 9 to determine the range of Cramér-Rao Bounds associated with our experiment.
D.4 Magnification and Alignment Correction
Figure 13 shows the experimental coronagraph throughput as a function of the scanned source position. Comparing the measured throughput to the theoretical throughput curve of our system revealed a scaling $m$ and shift $a$ between the true source displacements $y_{e}$ relative to the diffraction limit and the displacements computed from our system specifications $y_{e}^{{}^{\prime}}=m(y_{e}-a)$. The scaling and shift terms arise from experimental error in magnification and alignment of the optical axis respectively. After accounting for these discrepancies, we see the throughput of curve of our transformed (scaled and shifted) coordinates align with theoretical predictions.
@@ D.4 Magnification and Alignment Correction
Appendix E Simulated MPLC Cross-Talk Matrix
In this section, we address the complex cross-talk matrix of the MPLC mode sorter in greater detail. Historically, literature in spatial mode sorting has placed emphasis on defining the cross-talk matrix in terms of the power leakage between modes (amplitude-only characterization). However, complete characterization of the cross-talk requires estimation of both the amplitude and phase. Direct experimental characterization of the complex cross-talk matrix is a critical to forthcoming high-performance mode-sorting applications such as coronagraphy and spatial mode filtering. In this section we study the cross-talk matrix of the MPLC used in our experimental design with numerical field propagation simulations.
The goal of an MPLC is to passively map a given set of input modes $\{{\psi}_{i}\in\mathbb{C}^{M}\}_{i=1}^{K}$ into a desired set of output modes $\{{\phi}_{i}\in\mathbb{C}^{M}\}_{i=1}^{K}$. For concreteness, let $\Psi,\Phi\in\mathbb{C}^{M\times K}$ be matrices whose columns are (vectorized) orthonormal input and output modes represented in the pixel basis. In general, the number of pixel degrees of freedom $M$ is much greater than the number of modes $K$ that one wishes to sort. Additionally, we take the modes to be orthonormal such that $\Psi^{\dagger}\Psi=\Phi^{\dagger}\Phi=I_{K}$. The cross-talk matrix $\Omega\in\mathbb{C}^{K\times K}$ describes the interferometric connection between $\Psi$ and $\Phi$ within the MPLC transfer matrix:
$T=\Phi\Omega^{\dagger}\Psi^{\dagger}$
Under this decomposition of the transfer matrix $T$, the action of the MPLC on an arbitrary incident field can be understood as follows: First, project the arriving field onto the basis of input modes via $\Psi^{\dagger}$. Then, interfere light in the input mode channels according to the cross-talk matrix $\Omega$. Finally, expand the mixed channels back into the pixel basis via $\Phi$. Note that if $\Omega=I_{K}$, then the MPLC perfectly maps the input modes to the output modes in a bijective fashion. In the most general case however, $\Omega$ admits a singular value decomposition which comprehensively handles interferometric mixing between mode channels as well as loss through the device.
Figure 14: (a) Amplitude and (b) phase of the complex cross-talk matrix calculated from numerical simulations. (c) Simulated target and output fields of the mode sorter at the sorting plane after propagating the Zernike modes through the MPLC. The entries of the cross-talk matrix are the pair-wise projections (overlap integrals) between target modes and output modes.
A simple way to calculate the elements of the complex cross-talk matrix in simulation is by forward propagating each input mode independently and taking projection between the resulting output field at the sorting plane and the target output mode. That is we perform the computation,
$\Omega_{ij}=({\phi}_{i}^{\dagger}T{\psi}_{j})^{*},\qquad T=H_{\tau}D_{\tau}\cdots H_{1}D_{1}$
where we compute the transfer matrix via alternating application of phase masks (diagonal matrices $D_{i}$) inter-spaced by plane-to-plane light propagation (unitary matrices $H_{i}$). Figure 14(a,b) shows the simulated cross-talk of our MPLC device. We note that the this simulated cross-talk matrix is not unitary as the forward propagated modes span a slightly different subspace compared to the target mode set. This is visually evident in Figure 14(c) where residual speckles in the output fields are observed to lie outside the designated regions of support for the target modes. Contrast this matrix to the experimental cross-talk matrix displayed in Figure 5 which we evaluated using the methods described in Appendix D(1). Importantly, for calibrating the experimental cross-talk matrix, we assumed loss was negligible and forced the requirement that $\Omega$ be unitary. Despite this reduction in the degrees of freedom, the unitary cross-talk matrix calibrated for our bench-top implementation of the coronagraph provides an experimental model that aligns reasonably well with our measured results.
Appendix F Numerical Comparisons to Other Coronagraphs
Figure 15: Comparison of CRLBs for unbiased estimators of the exoplanet position under measurements with different coronagraphs. Each coronagraph measurement model includes the same empirical background noise characterized in our experimental system.
@@ Appendix F Numerical Comparisons to Other Coronagraphs
$\mathbf{q}(\vec{r}_{0})=|U\Sigma V^{\dagger}\mathbf{u}(\vec{r}_{0})|^{2}$ (27)
where $U,V\in\mathbb{C}^{M\times M}$ are unitary matrices and $\Sigma\in\mathbb{R}^{M\times M}$ is a diagonal matrix of singular values. Figure 15 shows the Craḿer-Rao lower bounds associated with each coronagraph. We see that the idealized version of our experimental setup (Experiment 4-Mode) outperforms both the PIAACMC and Vortex coronagraph for over sub-diffraction exoplanet locations in the range $|y_{e}|<=0.4\sigma$ which corresponds to where the CFI of the truncated mode basis begins to depart from the QFI limit as shown in Figure 3(b).
Figure 16: Numerical evaluation of the MLE bias and associated CRLB curve under experimental model. This CRLB curve is equivalent to that used in Figure 9(b)
Appendix G Cramer-Rao Lower Bound for the Biased MLE
The maximum likelihood estimator is known to be an unbiased estimator in the asymptotic limit of many repeated measurements (here the term ’measurement’ formally refers to the detection of a photon). While this assumption is increasingly valid for larger star-planet separations, wherein the number of detected exoplanet photons is large, it becomes invalid for small star-planet separations as the vast majority of exoplanet photons are discarded by the coronagraph. This motivates the analysis of Equation 11 in the main text. The expectations are $\mathbb{E}_{\mathbf{Y}|y_{e}}[\cdot]$ are taken over the joint distribution on the measurement outcome $\mathbf{Y}$ given by,
$P(\mathbf{Y}|y_{e})$ $=\prod_{j=1}^{M}P(Y_{j}|y_{e})$ (28a)
$P(Y_{j}=n|y_{e})$ $=\frac{e^{-\tau_{j}(y_{e})}\tau_{j}(y_{e})^{n}}{n!}$ (28b)
$\tau_{j}(y_{e})$ $=\Lambda_{0}p_{j}(y_{e})+\Lambda_{B}p_{B_{j}}$ (28c)
Finally, $I(y_{e})$ is the total classical Fisher Information defined as,
$I(y_{e})=\sum_{\mathcal{D}}P(\mathbf{Y}|y_{e})\bigg{[}\frac{d}{dy_{e}}\log P(\mathbf{Y}|y_{e})\bigg{]}^{2}$ (29)
where $\mathcal{D}$ is the event space of possible measurement outcomes for $\mathbf{Y}$. In Figure 16(a) we plot the bias of the MLE which we numerically computed by Monte-Carlo sampling the probabilistic measurement model of equation 3 to determine the expectation of the MLE $\mathbb{E}_{\mathbf{Y}|y_{e}}[\check{y}_{e}(y_{e})]$. We see that near the optical axis, the MLE is clearly biased as it no longer operates in the asymptotic regime. In Figure 16(b) we show the corresponding biased CRLB for the MLE of the exoplanet position.